Four times Marie's age minus twice Lisa's age is -10. Lisa is three times as old as Marie. How old is Marie?
Answer by jorel1380(3719)

Let Marie be m. Then Lisa would be 3m. So:
4m-2(3m)=-10
-2m=-10
m=5
Marie is 5; Lisa is 15. ☺☺☺☺